PREDICATE LOGIC.(25 pts)

A. Let P(x) be the statement x 2 > x4. If the domain consists of the integers,

what are the truth values?

1. P(0)

2. P(-1)

3. P(1)

4. P(2)

5. ∃xP(x)

6. ∀xP(x)


1
Expert's answer
2021-04-28T16:37:04-0400

Let "P(x)" be the statement "x^2 > x^4". If the domain consists of the integers, let us find the truth values:


1. Taking into account that it is not true that "0>0", we conclude that "P(0) =F."

2. Since "(-1)^2 =1= (-1)^4", we have that "P(-1) =F."

3. Taking into account that "1^2 =1= 1^4", we have that "P(1) =F."

4. Since "2^2=4<16=2^4", we conclude that "P(2) =F."

5. Taking into account that "x^2\le x^4" for any integer "x", we conclude that "\u2203xP(x)=F."

6. Since "x^2\le x^4" for any integer "x", we conclude that "\u2200xP(x)=F."